# Lazy-DaSH: Lazy Approach for Hypergraph-based Multi-robot Task and Motion Planning

### v1

@@ IV-B2 Query (Motion-extended Hypergraph)
The motion-extended hypergraph is defined as $\mathcal{H}_{\texttt{ME}}=(\mathcal{V}_{\texttt{ME}},\mathcal{E}_{\texttt{ME}})$, similar to $\mathcal{H}_{\texttt{TE}}$ but with motion information. Each motion-extended vertex $v_{\texttt{{ME}}}=\langle v_{\mathcal{M}},\Pi_{v_{\texttt{{ME}}}^{\texttt{{source}}}v_{\texttt{{ME}}}}\rangle\in\mathcal{V}_{\texttt{ME}}$ consists of a motion vertex $v_{\mathcal{M}}\in\mathcal{V}_{\mathcal{M}}$ and a motion-extended transition history $\Pi_{v_{\texttt{{ME}}}^{\texttt{{source}}}v_{\texttt{{ME}}}}$ that stores a history extending from the motion-extended source vertex $v_{\texttt{{ME}}}^{\texttt{{source}}}=\langle v_{\mathcal{M}},\emptyset\rangle$ to $v_{\texttt{{ME}}}$. Each motion-extended hyperarc $E_{\texttt{ME}}=\langle\texttt{Tail},\texttt{Head},E_{\mathcal{M}}\rangle\in\mathcal{E}_{\texttt{ME}}$ includes the information about the tail, head, and motion hyperarc that contributes to the history transitions. The motion query process begins at $v_{\texttt{{ME}}}^{\texttt{{source}}}$, expands the motion hyperarcs guided by the task plan, and terminates at $v_{\texttt{{ME}}}^{\texttt{{sink}}}$. The resulting motion history $\Pi_{v_{\texttt{ME}}^{\texttt{{source}}}v_{\texttt{ME}}^{\texttt{{sink}}}}$ in $v_{\texttt{{ME}}}^{\texttt{{sink}}}$ provides a motion plan that represents an optimistic schedule. This schedule ensures collision-free paths within each composition space but does not consider potential collisions with other moving bodies. The conflict resolution layer resolves these issues by refining individual motion plans to generate a collision-free schedule for all moving entities.
IV-C Conflict Resolution Layer
@@
IV-C1 Conflicts
@@
There are three types of motion conflicts, arising from interactions between two motion vertices, two motion hyperarcs, or a motion vertex and a motion hyperarc. A vertex-vertex conflict occurs when two objects are in a colliding position and is often handled by resampling the object’s positions within the planning area. A hyperarc-hyperarc conflict involves a collision between two moving entities and is generally resolved by adjusting the timing of the motion using a motion scheduling algorithm, or by replanning with path constraints to ensure the movable bodies avoid each other. A hyperarc-vertex conflict can arise when a static object obstructs the moving entity and is addressed by replanning the motion with constraints to navigate around the obstruction or by reordering the task sequence for colliding entities to pass through in a coordinated order. It is worth noticing that motion conflicts can be detected both in the motion query phase in the motion planning layer and the conflict resolution layer (Fig. 2).
IV-C2 Constraint Feedback
An overview of the constraint feedback is illustrated in Fig. 2. The hierarchical structure, comprising the task planning layer and the motion planning layer, enables the resolution of conflicts within these two distinct levels. The constraints are applied backward through the hierarchy. The motion conflicts are first addressed at the motion planning level by replanning the conflicting motions to find alternative motions. If these attempts consistently fail (e.g., timeout), it is assumed that the current task plan cannot be executed with a feasible motion, leading to the generation of task constraints. These task constraints are designed to produce an alternative task sequence that seeks a feasible motion plan.
IV-D Discussion
@@
As discussed at the beginning of Section IV, both Lazy-DaSH and DaSH can be categorized as hybrid and interleaved planning but differ in their emphasis on sequencing and constraint satisfaction, where DaSH prioritizes constraint satisfaction and Lazy-DaSH prioritizes sequencing. The approach for Lazy-DaSH aims to alleviate the exponential growth in representation size by identifying the minimum constraints required.
@@
Comparing how DaSH and Lazy-DaSH manage their state space representations effectively illustrates the differences between the two approaches. From the perspective of the search process, the search space expands during the task/motion representation construction phases (blue arrows in Fig. 1) and is queried by the task/motion query phase (green and red arrows in Fig. 1). DaSH constructs the representations in two immediate sequential phases, exhaustively adding every feasible transition and motion for each robot, object, and their interactions into the representation. This approach leads to a continuous expansion of search space across the representation construction layers (blue arrows). Then the combined task and motion planning query phase finds a solution within the constructed representation (green arrows). Conversely, Lazy-DaSH in Fig. 1 scopes the search space by utilizing the task query phase (red arrows) immediately after constructing the task space representation (top blue arrows). This occurs before further expanding search space in the subsequent motion representation construction phase (bottom blue arrows). This approach results in a more compact search space compared to DaSH, containing only the essential information, which facilitates faster query processes.
(a) Sort: 4 robots
(b) Sort: 8 robots
(c) Wall: 5 robots 1 wall
(d) Wall: 5 robots 3 walls
(e) Shelf-Wall: 2 robots 2 shelves
(f) Shelf-Wall: 4 robots 4 shelves
(g) Lab
@@
Figure 3: Four different types of experiment scenarios. (a) and (b) show “Sort” scenarios where the initial clusters of objects are represented within colored circles, and each group must be moved to the matching square boxes. (c) and (d) represent “Wall” scenarios, featuring different numbers of walls. (e) and (f) illustrate the “Shelf-Wall” scenario, showing the start and goal locations of the blocks. Finally, (g) is the “Lab” scenario, involving 3-axis gantry robots along with descriptions of the problem entities.
@@
Figure 4: A hardware experiment for the Shelf-Wall scenario. The top panels of the shelves have been removed for better visibility. A full video of demonstration can be found at: https://youtu.be/bWQwy0dWAJI
Figure 5: Representation size comparison in terms of the number of vertex and hyperarc for Sort, Wall, Shelf-Wall, and Lab scenarios. The y-axis represents the number of vertex/hyperarc on a logarithmic scale, while the x-axis denotes the number of objects.
@@
Figure 6: A comparison of motion representation construction time (1), task and motion query time (2), conflict resolution time (3), and total planning time (1+2+3) is presented for the Sort, Wall, Shelf, and Lab scenarios. In DaSH, the query time reflects the combined task and motion planning time, whereas in Lazy-DaSH, the query time is the sum of the task query and motion query times. The y-axis represents planning time on a logarithmic scale and is kept consistent across each scenario to clearly compare the proportion of each time component (1, 2, and 3) in the total planning time (1+2+3).
@@ V Validation
This section details the validation of Lazy-DaSH. We evaluated Lazy-DaSH across four scenarios, demonstrating its scalability and efficient constraint management across the hierarchical structure. Lazy-DaSH achieves a more compact representation and significantly faster total planning times, demonstrating superior scalability with twice the number of robots and objects. Notably, Lazy-DaSH improves planning times by an order of magnitude compared to DaSH [2], which itself has already outperformed the best existing method, Synchronized Multi-Arm Rearrangement (SMART) [25], by up to three orders of magnitude.
@@ V-A Evaluation Criteria
As discussed in Section IV-D, Lazy-DaSH maintains a compact yet sufficient representation to ensure successful queries. This results in a smaller representation size compared to DaSH, leading to faster queries and improved scalability. To assess these advantages, we evaluate Lazy-DaSH in four distinct scenarios that highlight its capabilities, measuring representation size and overall planning time as the number of robots and objects increases, as shown in Figs. 5 and 6. A detailed analysis is discussed in Section V-D.
@@ V-B Method Descriptions
In this section, we provide implementation details of Lazy-DaSH and DaSH.
V-B1 DaSH
Probabilistic Roadmap (PRM) [3] is used as a motion planner and motion feasibility is all validated during the $\mathcal{H}_{\mathcal{M}}$ construction phase. This validation eliminates the need for further motion validation during both the query phase and the conflict resolution phase. A variant of CBS-MP [28] is employed as a motion scheduler to sequence configurations over time in the conflict resolution layer.
V-B2 Lazy-DaSH
Lazy-Probabilistic Roadmap (Lazy-PRM) [5] is used as the motion planner during the $\mathcal{H}_{\mathcal{M}}$ construction stage, with motion feasibility evaluated during the motion query and conflict resolution phases. The motion scheduler in the conflict resolution phase is identical to the one used in the DaSH method.
Both Lazy-DaSH and DaSH were implemented in C++ using the Parasol Planning Library. The experiments were conducted on a desktop computer equipped with an Intel Core i9-14900K CPU at 3.2 GHz and 64 GB of RAM.
@@ V-C Scenarios
Lazy-DaSH and DaSH are evaluated on the multi-manipulator object rearrangement problem, where manipulators are tasked to transport blocks from the start state to the goal state by performing grasp and hand-over actions. The cube-shaped blocks have randomly generated start and goal positions, ensuring that at least one robot can grasp them and transfer them to the goal.
We demonstrate four key capabilities: 1) efficient planning for large numbers of robots, 2) an effective constraint feedback mechanism for identifying infeasible robot interactions, 3) resolving geometric constraints, and 4) the ability to solve problems that require more than one step for task completion.
To evaluate these capabilities, we designed four scenarios: Sort, Wall, Shelf-Wall, and Lab, as illustrated in Fig. 3. In each scenario, we progressively increase task complexity by adding more robots, resulting in a greater number of interactions needed to transfer objects to their goal locations. Additionally, the number of objects gradually increases, further intensifying the task complexity. These increments expand the size of both the task space hypergraph and the motion hypergraph, significantly increasing the computational complexity of the query process.
All manipulators in the Sort, Wall, and Shelf-Wall scenarios are UR5e arms equipped with Hand-e grippers, and the Lab scenario uses gantry robots with customized end effectors tailored to task requirements. We demonstrate a hardware experiment for our Shelf-Wall scenario, which represents a more complex problem compared to the Shelf experiment demonstrated in DaSH [2]. The corresponding video is provided in the link shown in Fig.4.
@@
V-C1 Sort
To assess Lazy-DaSH’s efficiency in large-scale planning, we designed Sort scenarios where four to sixteen manipulators collaborate to transport colored objects to designated boxes as shown in Figs. 3(a) and 3(b). This is essentially the same as the “cross sort” case described in DaSH paper [2], where objects must be delivered to the opposite side of the workstation. As the number of robots increases, this setup requires more complex coordination of task and motion, as each object must be handed over multiple times among the manipulators. It is important to note that the minimum number of required interaction grows as the number of robots increases for transfering each object to the goal location. The results show that Lazy-DaSH significantly improves scalability, successfully handling up to 8 robots with 20 objects, whereas DaSH could only handle up to 4 robots with 12 objects.

### v2

@@ IV-D2 Query (Motion-extended Hypergraph)
The motion-extended hypergraph is defined as $\mathcal{H}_{\texttt{ME}}=(\mathcal{V}_{\texttt{ME}},\mathcal{E}_{\texttt{ME}})$, similar to $\mathcal{H}_{\texttt{TE}}$ but with motion information. Each motion-extended vertex $v_{\texttt{{ME}}}=\langle v_{\mathcal{M}},\Pi_{v_{\texttt{{ME}}}^{\texttt{{src}}}v_{\texttt{{ME}}}}\rangle\in\mathcal{V}_{\texttt{ME}}$ consists of a motion vertex $v_{\mathcal{M}}\in\mathcal{V}_{\mathcal{M}}$ and a motion-extended transition history $\Pi_{v_{\texttt{{ME}}}^{\texttt{{src}}}v_{\texttt{{ME}}}}$ that stores a history extending from the motion-extended source vertex $v_{\texttt{{ME}}}^{\texttt{{src}}}=\langle v_{\mathcal{M}},\emptyset\rangle$ to $v_{\texttt{{ME}}}$. Each motion-extended hyperarc $E_{\texttt{ME}}=\langle\texttt{Tail},\texttt{Head},E_{\mathcal{M}}\rangle\in\mathcal{E}_{\texttt{ME}}$ includes the information about the tail, head, and motion hyperarc that contributes to the history transitions. The motion query process begins at $v_{\texttt{{ME}}}^{\texttt{{src}}}$, expands the motion hyperarcs guided by the task plan, and terminates at $v_{\texttt{{ME}}}^{\texttt{{sink}}}$.
When querying motions along the unvalidated schedule, motion constraints are applied to prevent traversal through regions defined by the boundaries of colliding objects and robots. These regions are identified in the subsequent layer, specifically motion conflict resolution layer, and incorporated into replanning requests. The motion constraint is defined as $\mathcal{C}_{\mathcal{M}}=(\Pi_{v_{\texttt{TE}}^{\texttt{src}}v_{\texttt{TE}}^{\texttt{1}}},\Pi_{v_{\texttt{TE}}^{\texttt{src}}v_{\texttt{TE}}^{\texttt{2}}},p)$, which consists of the motion histories leading to the colliding hyperarcs $E_{\texttt{TE}}^{1}$ and $E_{\texttt{TE}}^{2}$, together with $p$, a collision region in the task space derived from the geometries of the colliding objects. Since collisions are history-dependent, we preserve the transition histories of colliding hyperarcs and use their last elements to identify the conflicting motions. This allows the planner to mark regions for avoidance in context, enabling replanning without discarding the entire plan.
The resulting motion history $\Pi_{v_{\texttt{ME}}^{\texttt{{src}}}v_{\texttt{ME}}^{\texttt{{sink}}}}$ in $v_{\texttt{{ME}}}^{\texttt{{sink}}}$ provides a motion plan that represents an optimistic schedule. This schedule ensures collision-free paths within each composition space but does not consider potential collisions with other moving bodies. The conflict resolution layer resolves these issues by refining individual motion plans to generate a collision-free schedule for all moving entities.
IV-E Motion Conflict Resolution Layer
@@
IV-E1 Motion Conflicts
There are two types of conflicts in the motion conflict resolution layer, arising from interactions between two motion hyperarcs, or between a motion vertex and a motion hyperarc. Note that conflicts between two motion vertices are already identified in the task conflict detection phase and resolved by replanning the task plan in the task planning layer (Section IV-B).
@@ IV-E1 Motion Conflicts
A hyperarc–hyperarc conflict involves a collision between two moving entities and is resolved by adjusting the timing of the motion using a motion scheduling algorithm, or by replanning with path constraints to ensure the moving bodies avoid each other. A hyperarc–vertex conflict arises when a static object obstructs a moving entity and is addressed either by replanning the motion with constraints to navigate around the obstruction using motion constraints ($\mathcal{C}_{\mathcal{M}}$), or by reordering the task sequence so that the colliding entities pass through in a coordinated order using task constraints ($\mathcal{C}_{\mathcal{T}}$), and it is further described in the subsequent section.
IV-E2 Motion Conflict Resolution and Constraint Feedback
Motion conflicts can be detected and addressed in both the motion query phase of the motion planning layer and the motion conflict resolution layer (Fig. 2), but the replanning strategies differ. In the motion query phase, failures are handled by searching for alternative motions or refining the roadmap. In contrast, the motion conflict resolution layer formulates a subproblem that defines a new composite space, enabling replanning in a larger motion search space. This approach is inspired by the Adaptive Robot Coordination (ARC) method [46] and adapted to the task and motion planning framework, where task dependencies coordinate the initiation and termination of motions to ensure alignment with the timing of related tasks, as shown in Algorithm 3 and Figure 4.
Algorithm 3 Scheduled Adaptive Robot Coordination
1:Task space hypergraph $\mathcal{H}_{\mathcal{T}}$, motion hypergraph $\mathcal{H}_{\mathcal{M}}$, task transition history $\Pi_{v_{\texttt{TE}}^{\texttt{{src}}}v_{\texttt{TE}}^{\texttt{{sink}}}}$, motion transition history $\Pi_{v_{\texttt{ME}}^{\texttt{{src}}}v_{\texttt{ME}}^{\texttt{{sink}}}}$.
[…]
3: // Build Dependency Graph
[…]
5: // Collect Mapping $E_{\mathcal{T}}\rightarrow E_{\mathcal{M}}$
[…]
7: // Find Motion Conflicts
[…]
9:while $C_{\mathcal{M}}\neq\emptyset$
10:  $\mathcal{Q},G\leftarrow\textsc{CreateSubProblem}(C_{\mathcal{M}},M,D)$
11:  $\mathcal{E}_{\mathcal{M}}\leftarrow\textsc{SolveSubProblem}(M,\mathcal{Q})$
12:  if $\mathcal{E}_{\mathcal{M}}\neq\emptyset$ $\triangleright$ conflict resolved
13:   $\textsc{UpdateMotionHypergraph}(\mathcal{H}_{\mathcal{M}},\mathcal{E}_{\mathcal{M}}^{\prime})$
14:   $C_{\mathcal{M}}\leftarrow\textsc{FindMotionConflict}(M,D)$
15:  else
16:   return $\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}}$
Figure 4: Illustration of scheduled adaptive robot coordination.
IV-F Discussion
This section compares Lazy-DaSH to DaSH by examining their emphasis on sequencing versus constraint satisfaction, their approaches to managing state space representations, and provides a discussion of probabilistic completeness.
IV-F1 Comparison with DaSH
@@
As discussed at the beginning of Section IV-A, both Lazy-DaSH and DaSH can be categorized as hybrid and interleaved planning but differ in their emphasis on sequencing and constraint satisfaction, where DaSH prioritizes constraint satisfaction and Lazy-DaSH prioritizes sequencing. The approach for Lazy-DaSH aims to alleviate the exponential growth in representation size by identifying the minimum constraints required.
@@
Comparing how DaSH and Lazy-DaSH manage their state space representations effectively illustrates the differences between the two approaches. From the perspective of the search process, the search space expands during the task/motion representation construction phases (blue arrows in Fig. 1) and is queried by the task/motion query phase (green and red arrows in Fig. 1). DaSH constructs the representations in two immediate sequential phases, exhaustively adding every feasible transition and motion for each robot, object, and their interactions into the representation. This approach leads to a continuous expansion of search space across the representation construction layers (blue arrows). Then the combined task and motion planning query phase finds a solution within the constructed representation (green arrows). Conversely, Lazy-DaSH in Fig. 1 narrows the search space by invoking the task query phase (red arrows in Task Query) immediately after constructing the task space representation (blue and red arrows in Task Space Construction). Furthermore, the task conflict detection layer postpones motion space expansion until a valid task plan is found (red arrows in Task Conflict Detection) and expands the task space elements only when necessary. This occurs before further expanding search space in the subsequent motion representation construction phase (blue arrows in Lazy Motion Space Construction phase). This approach results in a more compact search space compared to DaSH, containing only the essential information, which facilitates faster query processes.
IV-F2 Theoretical Properties
This section elaborates on the probabilistic completeness of Lazy-DaSH. The property arises from its hierarchical, iterative structure, where representation construction and constraint-driven expansion ensure that the search space grows whenever the current representation is insufficient.
The guarantee builds on the properties of the task planning and motion planning components. For a given task-extended hypergraph $\mathcal{H}_{\texttt{TE}}$, a depth-first search with backtracking is complete and will always find a task plan if one exists within the representation $\mathcal{H}_{\mathcal{T}}$. For the motion layer, Lazy-PRM is probabilistically complete, so if a feasible motion exists, the probability of finding it approaches one as the roadmap becomes denser, with unvalidated edges eventually being evaluated.
[…]
The overall process follows a consistent cycle in which the planner queries on the current $(\mathcal{H}_{\mathcal{T}},\mathcal{H}_{\mathcal{M}})$, detects failure, generates constraints, expands the representations, and then re-queries. Two conditions ensure that this loop provides probabilistic completeness. First, expansions are monotone, meaning they only add task and motion elements and never permanently remove a potentially feasible plan. Second, expansions are exhaustive, meaning that as iteration proceeds, every feasible task sequence or motion configuration has a nonzero probability of being generated, and with increasing budgets the probability that all relevant interactions are eventually attempted approaches one.
Because query failures always trigger constructive expansion, Lazy-DaSH systematically enlarges its search space until it contains a feasible solution. Therefore, under the standard assumptions for Lazy-PRM together with monotone and exhaustive expansion of $\mathcal{H}_{\mathcal{T}}$ and $\mathcal{H}_{\mathcal{M}}$, Lazy-DaSH is probabilistically complete.
(a) Sorting: 4 robots
(b) Sorting: 16 robots
(c) Wall: 8 robots / 2 wall
(d) Wall: 8 robots / 6 walls
(e) Shelf-wall: 2 robots / 2 shelves
(f) Shelf-wall: 4 robots / 4 shelves
(g) Stocking: 2 robots / 3 and 2 layers
(h) Lab
@@ IV-F2 Theoretical Properties
Figure 5: Five different types of experiment scenarios. (a) and (b) show “Sorting” scenarios where the initial clusters of objects are represented within colored circles, and each group must be moved to the matching square boxes. (c) and (d) represent “Wall” scenarios, featuring different numbers of walls. (e) and (f) illustrate the “Shelf-wall” scenario, showing the start and goal locations of the blocks. Finally, (g) is the “Lab” scenario, involving 3-axis gantry robots along with descriptions of the problem entities.
@@
Figure 6: A hardware experiment for the Shelf-wall scenario. The top panels of the shelves have been removed for better visibility.
Figure 7: Comparison of task and motion representation sizes between Lazy-DaSH and the original framework, DaSH, measured by the number of vertices and hyperarcs in the Sorting, Wall, Shelf-wall, Stocking, and Lab scenarios. The y-axis shows the number of vertices/hyperarcs on a logarithmic scale, and the x-axis indicates the number of objects.
@@
Figure 8: A comparison of motion representation construction time (1), task and motion query time (2), conflict resolution time (3), and total planning time (1+2+3) is presented for the Sorting, Wall, Shelf-wall, Stocking, and Lab scenarios. In DaSH, representation time corresponds to motion construction and query time to the combined task–motion planning. In Lazy-DaSH, by contrast, task- and motion-level times are separated, with dashed lines indicating task and solid lines indicating motion. The y-axis represents planning time on a logarithmic scale and is kept consistent across each scenario to clearly compare the proportion of each time component (1, 2, and 3) in the total planning time (1+2+3).
@@ V Validation
This section details the validation of Lazy-DaSH. We evaluated Lazy-DaSH across five scenarios, demonstrating its scalability and efficient constraint management across the hierarchical structure. Lazy-DaSH achieves a more compact representation and significantly faster total planning times, demonstrating superior scalability with twice the number of robots and objects than DaSH [2]. Notably, Lazy-DaSH achieves planning times that are up to an order of magnitude faster than the original framework, which itself has already demonstrated up to three orders of magnitude improvement over the coupled and synchronous planner, Synchronized Multi-Arm Rearrangement (SMART)  [34].
@@ V-A Evaluation Criteria
As discussed in Section IV-A, both Lazy-DaSH and DaSH utilize a hypergraph-based representation to efficiently capture the multi-manipulator object rearrangement problem. This efficiency arises from the hybrid approach, in contrast to graph-based methods that represent the composite state space. The hypergraph representation encodes the problem compactly, and querying over this hybrid representation enables faster performance compared to composite state space representations. Compared to DaSH, Lazy-DaSH preserves this advantage while further improving representation management through its constraint feedback mechanism.
The benefit of the hybrid approach enabled by the hypergraph-based representation has already been validated in the original DaSH work [2] where it demonstrated up to three orders of magnitude speed up in planning times over state-of-the-art composite approaches. Here we compare directly against DaSH to evaluate the impact of the contributions presented in this paper. We show that Lazy-DaSH further enhances the hybrid representation by achieving more compact representations and faster query times in complex scenarios. To assess these advantages, we measure both representation size and overall planning time as the complexity of each problem increases, comparing results with DaSH (Figs. 7 and 8). A detailed analysis is provided in Section V-D.
@@ V-B Method Descriptions
In this section, we provide implementation comparison of Lazy-DaSH and DaSH.
For motion planner, DaSH employs the Probabilistic Roadmap (PRM) method [3], whereas Lazy-DaSH employs the Lazy Probabilistic Roadmap (Lazy-PRM) method [5]. For a fair comparison of the query process, both methods use the scheduled adaptive robot coordination strategy introduced in Section IV-E for motion conflict resolution, but they differ in their feedback trigger criteria. DaSH follows the CBS-MP framework with probabilistic break rules [47], which determine whether to continue expanding the constraint tree or to expand the roadmaps before restarting the query. Both Lazy-DaSH and DaSH were implemented in C++, and the experiments were conducted on a desktop computer equipped with an Intel Core i9-14900K CPU at 3.2 GHz and 64 GB of RAM.
@@ V-C Scenarios
Lazy-DaSH and DaSH are evaluated on the multi-manipulator object rearrangement problem, where manipulators are tasked to transport blocks from the start state to the goal state by performing grasp and hand-over actions. The cube-shaped blocks have randomly generated start and goal positions, ensuring that at least one robot can grasp them and transfer them to the goal. For grasp poses, we consider the sides, top, and bottom as possible options.
We demonstrate five key capabilities: (1) scalable planning for large-scale multi-robot systems, (2) an effective constraint feedback mechanism for identifying infeasible robot interactions, (3) resolution of geometric constraints, (4) expansion of the task search space to handle non-monotonicity, and (5) the ability to solve tasks requiring multiple steps for completion.
To evaluate these capabilities, we designed five scenarios: Sorting, Wall, Shelf-wall, Stocking, and Lab, as illustrated in Fig. 5. In each scenario, we progressively increase task complexity based on the features we aim to demonstrate. In all cases, task complexity grows with the number of objects, and the robots must coordinate and interact to complete the tasks. These increments expand the size of both the task space hypergraph and the motion hypergraph, significantly increasing the computational complexity of the query process.
All manipulators in the Sorting, Wall, Shelf-wall, and Stocking scenarios are UR5e arms equipped with Hand-e grippers, and the Lab scenario uses gantry robots with customized end effectors tailored to task requirements. We demonstrate a hardware experiment for our Shelf-wall scenario, which represents a more complex problem compared to the Shelf experiment demonstrated in DaSH [2]. The corresponding video is provided in the link shown in Fig.6.
@@
V-C1 Sorting
To assess Lazy-DaSH’s efficiency in large-scale planning, we designed Sorting scenarios where four to sixteen manipulators collaborate to transport colored objects to designated boxes as shown in Figs. 5(a) and 5(b). This is essentially the same as the “cross sorting” case described in DaSH paper [2], where objects must be delivered to the opposite side of the workstation. As the number of robots increases, this setup requires more complex coordination of task and motion, as each object must be handed over multiple times among the manipulators. It is important to note that the minimum number of required interaction grows as the number of robots increases for transfering each object to the goal location. The results show that Lazy-DaSH significantly improves scalability, successfully handling 16 robots with more than 60 objects, whereas DaSH could only handle up to 4 robots with 14 objects.